OpenStudy (anonymous):

What is the value of (9^7)^3/14, in simplest terms?

OpenStudy (anonymous):

well first your going to have to narrow it down like always use paraentheses first. then do the other stuff and then narrow it down to is simplest form.

OpenStudy (solomonzelman):

\(\Large\color{black}{ \displaystyle \left( 9^7 \right)^{3/14}~~~\Rightarrow~~~9^{7 \times \frac{3}{14}} }\)

OpenStudy (ribhu):